OpenStudy (anonymous):

x^5-12x^3+32x=0 Can someone show me how they got this answer with steps because I need to know for other problems

OpenStudy (abb0t):

what are you being asked to do???

OpenStudy (triciaal):

is this answer to a problem or do you need to factor? if you need to factor first factor the x to get x(x^4 - 12 x^2 +32) 2 ? what is the sqrt (x^4)? and what are the factors of 32 that add up to -12? sqrt of x^4 = x^2 factors of 32 to use -8 and -4 x(x^2 -8)(x^2-4) x^2 -4 can be factored as the difference of 2 squares do you recall a^2 - b^2 = (a +b)(a-b) x^2-4 =(x+2)(x-2) now you have x(x^2-8)(x+2)(x-2)=0

OpenStudy (anonymous):

This is Calculus?